John plans to buy a vacation home in 10 years from now and wants to have saved $70,778 for a down payment. How much money should he place today in a saving account that earns 9.08 percent per year (compounded daily) to accumulate money for his down payment? Round the answer to two decimal places
Answer #1

Number of days in an year = 365

Number of days in 10 years = n = 10*365 = 3650

Daily Interest Rate = r = 0.0908/365

Amount deposited now = PV

Future Value required = FV = $70778

Hence, FV = PV(1+r)n

=> PV = FV/(1+r)n

= 70778/(1+0.0908/365)3650

= $28,550.12